1) Convert the following polar equation of r=2sintheta into a rectangular form.

2) List and find all of the sixth roots of 4.
Answer #1

sel?-0) Given that the polar equation na 2 Sino Convert the polar equation into rectangular form. Noul ours a El e ng sapts ysol? (2) The sixth roof of 4 are all solution- :326-4=0 (23)?-12-0 (x-2)(x+2) = 0 use thee identity (a b) = (a-b) (afb) (323-And the zeros of +3+372)3 = use the identity sum of cube x²+63 = (a+b) (a² a 6+62) 23+0) = (x+27)(x?- *.17 +°)=0 Hence the ze